# LLM Task Interference: An Initial Study on the Impact of Task-Switch in Conversational History

### v1

@@ Abstract
With the recent emergence of powerful instruction-tuned large language models (LLMs), various helpful conversational Artificial Intelligence (AI) systems have been deployed across many applications. When prompted by users, these AI systems successfully perform a wide range of tasks as part of a conversation. To provide some sort of memory and context, such approaches typically condition their output on the entire conversational history. Although this sensitivity to the conversational history can often lead to improved performance on subsequent tasks, we find that performance can in fact also be negatively impacted, if there is a task-switch. To the best of our knowledge, our work makes the first attempt to formalize the study of such vulnerabilities and interference of tasks in conversational LLMs caused by task-switches in the conversational history. Our experiments across 5 datasets with 15 task switches using popular LLMs reveal that many of the task-switches can lead to significant performance degradation.  [^1] Code available on GitHub.
@@
[…]
@@ 1 Introduction
The performance of LLMs is further boosted by leveraging in-context examples or few-shot examples of a particular task Brown et al. (2020); Smith et al. (2022); Thoppilan et al. (2022). In-context learning, by utilizing examples within the conversation history, enables LLMs to generate responses that are relevant and tailored to the contextual conversation. The auto-regressive nature of popular instruction-tuned (LLMs) suggests that the LLM generated response is conditioned on the entire conversation history. This underscores the sequential dependency and contextual awareness embedded within these models. While prompt sensitivity has been exploited by in-context learning to improve downstream performance, this sensitivity has also opened the door to vulnerabilities, where malicious actors can exploit prompt sensitivity for adverse purposes Greshake et al. (2023); Liu et al. (2023); Jiang et al. (2023b); Xu et al. (2023).
@@
  • We measure the task-switch sensitivity for popular LLMs of different sizes, where we observe that very large (175B) and small (7B) LLMs can both be susceptible to performance degradation from task-switch.
@@ 2 Related Work
Large Language Models (LLMs) are becoming a crucial building block of conversation-based virtual assistants OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023a); Anil et al. (2023). Leveraging in-context or few-shot examples, LLMs have demonstrated remarkable capabilities for downstream tasks  Brown et al. (2020). In contrast to the resource-intensive fine-tuning process Gao et al. (2020), in-context learning eliminates the need for parameter updates, while achieving state-of-the-art performance Rae et al. (2021); Smith et al. (2022); Thoppilan et al. (2022); Von Oswald et al. (2023); Chan et al. (2022); Akyürek et al. (2022); Hahn and Goyal (2023). However, despite its advantages, in-context learning tends to suffer from sensitivity to prompts, input distribution, and formats, which can potentially impact the model’s performance Liu et al. (2021); Zhao et al. (2021); Lu et al. (2021); Min et al. (2022); Liu and Wang (2023); Chang and Jia (2023).  Chang and Jia (2023) observe that the in-context examples implicitly bias the model. In our work, we aim to study the bias that arises due to chat history (in-context examples) when a user switches the task. Furthermore, recent works Liu et al. (2023); Greshake et al. (2023) have looked at the vulnerability of LLM to prompt injections and adversarial attacks. Unlike prompt injection, where a malicious prompt may be added to the conversation of LLM, our setting, is concerned with non-malicious task-switches. While a few recent works have investigated the reliance on shortcuts in conversation history Tang et al. (2023); Si et al. (2022); Weston and Sukhbaatar (2023), our work aims to evaluate prompt history sensitivity for a new task. Our work is also differentiated from the study topic change in Task-oriented Dialogue systems Xie et al. (2021); Xu et al. (2021); Yang et al. (2022) as we consider a stronger shift of task-switch from open dialogue LLMs.
@@ 3 Conversational Task-Switch
$r_{L+1}=\operatorname*{arg\,max}_{r}P_{\theta}(r|u_{L+1},\mathbf{h})$ (1)
In this work, we consider conversations with a single task-switch, where all user requests in the conversation history $\mathbf{h}$ belong to the same task and the final user request $u_{L+1}$ is a different task. We refer to the task associated with $\mathbf{h}$ as the conversation history task (CH task) $T_{h}$ where $\mathbf{h}\in T_{h}$ and the switched task associated with the final user request $u_{L+1}$ as the target task $T_{t}$ where $u_{L+1}\in T_{t}$.
When the tasks $T_{h}$ and $T_{t}$ are sufficiently different (as per human understanding of language and tasks), the conversation history $\mathbf{h}$ ideally must not impact the response, $r_{L+1}$. For a model robust to such a task-switches, $T_{h}\to T_{t}$, its response $r_{L+1}$ is conditionally independent of the conversation history,
@@
However, in practice, models can be sensitive to the conversation history, $\mathbf{h}$, which can harm the quality of the response $r_{L+1}$ after a task-switch, $T_{h}\to T_{t}$. We define $\tau(\cdot)$, the task-switch sensitivity of a model $\theta$, to measure the extent of this vulnerability. [^2] Theoretical and empirical implications of other definitions for task-switch sensitivity in Appendix E
@@
Task-switch sensitivity can be interpreted as:
@@ 4.1 Experimental Setup
We evaluate five different datasets covering a range of tasks: Gigaword (Graff et al., 2003); abstract algebra subset of Measuring Massive Multitask Language Understanding (MMLU; Hendrycks et al. (2021)), named MMLU AA; TweetQA (Xiong et al., 2019); Rotten Tomatoes (RT; Pang and Lee (2005)); and human-aging subset from the MMLU dataset (MMLU HA) in the Appendix.
@@
We explore the task-switch sensitivity of four popular models. We consider two open-source small models, Llama-7b-chat Touvron et al. (2023) and Mistral-7b-chat Jiang et al. (2023a); and two larger closed models, GPT-3.5 Brown et al. (2020) and GPT-4 OpenAI (2023). Zero-shot, absolute model performances are presented in Appendix B.
@@ 4.2 Results
In addition to the task-switch sensitivity $\tau(\cdot)$, we assess performance changes between the predictions in the presence of history and task-switch vs zero-shot. Table 2 and Table 3 showcases the impact of conversational task-switch with MMLU AA and Rotten Tomatoes as the target tasks, $T_{t}$ respectively [^3] The impact of task-switch for other datasets as the target tasks is given in Appendix C.1. As would be expected with in-context examples, the performance change in accuracy is generally positive. The negative trend for change in accuracy from $T_{h}\to T_{t}$, suggests that the task-switch causes performance degradation. For example, in the Gigaword summarization task as $T_{h}$ and MMLU AA as $T_{t}$, most models (GPT-3.5, Llama-7B and Mistral-7B) see a performance drop. Interestingly, for some models, the task-switch may increase performance; most prominently for Mistral-7B with Rotten Tomatoes as $T_{h}$ and MMLU AA as $T_{t}$.
@@
MMLU AA GPT-3.5 19.35 ***
GPT-4 8.62 ***
Llama-7B 3.57 31.51
Mistral-7B 28.57 1.12
Gigaword GPT-3.5 $-3.13$**
GPT-4 $0.00$**
Llama-7B $-10.71$ 5.23
Mistral-7B $-19.05$ 3.13
Rotten Tomatoes GPT-3.5 $6.45$**
GPT-4 $-13.11$**
Llama-7B $-15.63$ 9.91
Mistral-7B $18.18$ $0.83$
TweetQA GPT-3.5 $6.45$**
GPT-4 $-3.39$**
Llama-7B $0.00$ 6.37
Mistral-7B $-4.76$ 2.78
Table 2: Task-switch impact from CH-tasks ($T_{h}$) to target ($T_{t}$): MMLU AA and conversation length L = 6. Sensitivity not calculable for ***.
@@
Rotten Tomatoes GPT-3.5 3.00 ***
GPT-4 1.74 ***
Llama-7B 1.82 4.02
Mistral-7B 3.79 2.65
Gigaword GPT-3.5 $0.11$**
GPT-4 $-0.98$**
Llama-7B $1.82$ 1.98
Mistral-7B $-1.30$ 3.04
MMLU AA GPT-3.5 $-0.22$**
GPT-4 $0.76$**
Llama-7B $-5.69$ 3.37
Mistral-7B $0.97$ 1.39
TweetQA GPT-3.5 $-0.33$**
GPT-4 $-0.98$**
Llama-7B $3.76$ 2.77
Mistral-7B $-0.87$ 3.01
@@
Table 3: Task-switch impact from CH-tasks ($T_{h}$) to target ($T_{t}$): Rotten Tomatoes and conversation length L = 6. Sensitivity not calculable for ***.
@@
The sensitivity of different models to different task-switches can be compared fairly using the task-switch metric, $\tau(\cdot)$ The larger the value of $\tau(\cdot)$, the greater a model’s sensitivity to a specific task-switch. In Table 2 and Table 3, Llama-7B usually has the highest sensitivity to task-switches with for example $\tau=3.37$ for a switch from MMLU AA to Rotten Tomatoes and $\tau=9.91$ for task-switch from Rotten Tomatoes to MMLU AA. We observe a general trend between the change in accuracy and $\tau(\cdot)$ for task-switch scenarios for $T_{t}=$ Rotten Tomatoes where a negative change in performance also suggests very high task-switch sensitivity. In Figure 2, we plot the change in performance with increasing $T_{h}$ examples for MMLU AA dataset. Here we can clearly observe that in-context examples improve the predictive performance. Notably, the accuracy variation is more pronounced in smaller 7B models, likely due to their lower baseline performance, which is substantially improved by in-context learning. Performance fluctuations for conversation history, $\mathbf{h}$, can stem from two primary factors: a significant drop in the predicted probability for the zero-shot response, $r^{*}$, or a notable increase in the probability for an alternative response, $r$. The latter can result in substantial performance change while maintaining low sensitivity, $\tau(\cdot)$. By analyzing both performance changes and task-switch sensitivity, we gain deeper insights into the models’ adaptability to task-switches and the underlying dynamics influencing these shifts.
@@ 5 Conclusions and Future Work
This work formalizes and performs an initial investigation into the sensitivity of large language models (LLMs) to task-switch scenarios within conversational contexts. We introduce a task-sensitivity metric that can explain a model’s behavior to task-switches along with the performance change. By experimenting with various task-switch settings, we observe that even advanced models like GPT-4 exhibit vulnerabilities to task-switches. Our work additionally lays the foundation for future work on ‘side-channel’ vulnerabilities of LLMs to undesired information leakage/bias from the conversation history. Further work will focus on developing adaptive context management strategies within LLMs to mitigate the risk of task-switch sensitivity.
@@ 8 Acknowledgements
The appendix is structured as follows: Appendix A gives more details about the datasets, Appendix B reports the zero-shot absolute performance of all models on all tasks, Appendix C presents an ablation study on the conversation history length, Appendix D discusses the prompt templates used for each dataset, and Appendix E discusses other definitions for task-switch sensitivity.
@@ Appendix B Absolute Performance
When evaluating the target task with a conversation history, it is useful to compare the performance against a baseline with no conversation history ($\mathbf{h}=\text{Ø},L=0$). This is equivalent to evaluating in a zero-shot setting. This section reports the zero-shot performance for all the target task ($T_{t}$) datasets: MMLU HA in Table  5, MMLU AA in Table 6, RT in Table 7, Gigaword in Table  8 and TweetQA in Table 9. Also note that for the classification tasks (MMLU HA, MMLU AA, RT), we also report the number of responses for which we were unable to extract the answer (# Format Errors), which is further discussed in Appendix D. We evaluate on the test set with four LLMs (GPT-3.5, GPT-4, Mistral-7B, Llama-7B), which were all set to Temperature 0 for reproducability.

### v2

@@ Abstract
With the recent emergence of powerful instruction-tuned large language models (LLMs), various helpful conversational Artificial Intelligence (AI) systems have been deployed across many applications. When prompted by users, these AI systems successfully perform various tasks as part of a conversation. Such approaches typically condition their output on the entire conversational history to provide some sort of memory and context. Although this sensitivity to the conversational history can often lead to improved performance on subsequent tasks, we find that performance can in fact also be negatively impacted, if there is a task-switch. To the best of our knowledge, our work makes the first attempt to formalize the study of such vulnerabilities and interference of tasks in conversational LLMs caused by task-switches in the conversational history. Our experiments across 5 datasets with 15 task switches using popular LLMs reveal that many of the task-switches can lead to significant performance degradation. [^1] Code available on GitHub.
@@
Akash Gupta∗1  Ivaxi Sheth∗2  Vyas Raina1  Mark Gales1  Mario Fritz2 1University of Cambridge,  2CISPA Helmholtz Center for Information Security ag2118@cantab.ac.uk, {vr313, mjfg}@cam.ac.uk,  {ivaxi.sheth, fritz}@cispa.de
@@ 1 Introduction
The performance of LLMs is further boosted by leveraging in-context examples or few-shot examples of a particular task Brown et al. (2020); Smith et al. (2022); Thoppilan et al. (2022). In-context learning, by utilizing examples within the conversation history, enables LLMs to generate responses that are relevant and tailored to the contextual conversation. The auto-regressive nature of popular instruction-tuned (LLMs) suggests that the LLM-generated response is conditioned on the entire conversation history. This underscores the sequential dependency and contextual awareness embedded within these models. While prompt sensitivity has been exploited by in-context learning to improve downstream performance, this sensitivity has also opened the door to vulnerabilities, where malicious actors can exploit prompt sensitivity for adverse purposes Greshake et al. (2023); Liu et al. (2023); Jiang et al. (2023b); Xu et al. (2023).
@@
  • We measure the task-switch sensitivity for popular LLMs of different sizes, where we observe that LLMs of different sizes (7B to 175B) could be susceptible to performance degradation from the task-switch. [^2] Note that in some rare cases, the task-switch can lead to marginal performance improvements.
@@ 2 Related Work
Large Language Models (LLMs) are becoming a crucial building block of conversation-based virtual assistants OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023a); Anil et al. (2023). Leveraging in-context or few-shot examples, LLMs have demonstrated remarkable capabilities for downstream tasks  Brown et al. (2020). In contrast to the resource-intensive fine-tuning process Gao et al. (2020), in-context learning eliminates the need for parameter updates, while achieving state-of-the-art performance Rae et al. (2021); Smith et al. (2022); Thoppilan et al. (2022); Von Oswald et al. (2023); Chan et al. (2022); Akyürek et al. (2022); Hahn and Goyal (2023). However, despite its advantages, in-context learning tends to suffer from sensitivity to prompts, input distribution, and formats, which can potentially impact the model’s performance Liu et al. (2021); Zhao et al. (2021); Lu et al. (2021); Min et al. (2022); Liu and Wang (2023); Chang and Jia (2023).  Chang and Jia (2023) observe that the in-context examples implicitly bias the model. In our work, we aim to study the bias that may arise due to chat history (in-context examples) when a user switches the task. Furthermore, recent works Liu et al. (2023); Greshake et al. (2023) have looked at the vulnerability of LLM to prompt injections and adversarial attacks. Unlike prompt injection, where a malicious prompt may be added to the conversation of LLM, our setting, is concerned with non-malicious task-switches. While a few recent works have investigated the reliance on shortcuts in conversation history Tang et al. (2023); Si et al. (2022); Weston and Sukhbaatar (2023), our work aims to evaluate the influence of the conversation on a new task. Our work is also differentiated from the study topic change in Task-oriented Dialogue systems Xie et al. (2021); Xu et al. (2021); Yang et al. (2022) as we consider a stronger shift of task-switch from open dialogue LLMs.
@@ 3 Conversational Task-Switch
$r_{L+1}=\operatorname*{arg\,max}_{r}P_{\theta}(r|u_{L+1},\mathbf{h}).$ (1)
In this work, we consider conversations with a single task-switch, where all user requests in the conversation history $\mathbf{h}$ belong to the same task, and the final user request $u_{L+1}$ is a different task. We refer to the task associated with $\mathbf{h}$ as the conversation history task (CH task) $T_{h}$ where $\mathbf{h}\in T_{h}$ and the switched task associated with the final user request $u_{L+1}$ as the target task $T_{t}$ where $u_{L+1}\in T_{t}$.
When the tasks $T_{h}$ and $T_{t}$ are sufficiently different (as per human understanding of language and tasks), the conversation history $\mathbf{h}$ ideally must not impact the response, $r_{L+1}$. For a model robust to such task-switches, $T_{h}\to T_{t}$, its response $r_{L+1}$ is conditionally independent of the conversation history,
@@
However, in practice, models can be sensitive to the conversation history, $\mathbf{h}$, which can harm the quality of the response $r_{L+1}$ after a task-switch, $T_{h}\to T_{t}$. We define $\tau(\cdot)$, as a reference-free metric for the task-switch sensitivity of a model $\theta$, to measure the extent of this vulnerability. Theoretical and empirical implications of other definitions for task-switch sensitivity are considered in Appendix E.
@@
The task-switch sensitivity metric can be interpreted as:
@@ 4.1 Experimental Setup
We evaluate five different datasets covering a range of tasks: Gigaword (Graff et al., 2003); abstract algebra subset of Measuring Massive Multitask Language Understanding (MMLU; Hendrycks et al. (2021)), named MMLU AA; TweetQA (Xiong et al., 2019); Rotten Tomatoes (RT; Pang and Lee (2005)); and human-aging subset from the MMLU dataset (MMLU HA) in the Appendix I.
@@
We explore the task-switch sensitivity of four popular models. We consider two open-source small models, Llama2-7b-chat Touvron et al. (2023) and Mistral-7b-chat Jiang et al. (2023a); and two larger closed models, GPT-3.5 Brown et al. (2020) and GPT-4 OpenAI (2023). Zero-shot, absolute model performances are presented in Appendix B.
@@ 4.2 Results
We assess performance changes between the predictions in the presence of history and task-switch vs zero-shot. Table 2 and Table 3 showcases the impact of conversational task-switch with MMLU AA and Rotten Tomatoes as the target tasks, $T_{t}$ respectively [^3] The impact of task-switch for other datasets as the target tasks is given in Appendix C.1. As would be expected with in-context examples, the performance change in accuracy is generally positive. The negative trend for change in accuracy from $T_{h}\to T_{t}$, suggests that the task-switch causes performance degradation. For example, in the Gigaword summarization task as $T_{h}$ and MMLU AA as $T_{t}$, most models (GPT-3.5, Llama-7B and Mistral-7B) see a performance drop. Interestingly, for some models, the task-switch may increase performance; most prominently for Mistral-7B with Rotten Tomatoes as $T_{h}$ and MMLU AA as $T_{t}$.
@@
MMLU AA GPT-3.5 $\ \ \ 16.19_{\pm 5.51}$
GPT-4 $\ \ \ \ 0.26_{\pm 3.12}$
Llama-7B $\ \ \ -1.41_{\pm 14.93}$ $31.51$
Mistral-7B $\ \ \ \ 32.91_{\pm 12.08}$ $1.12$
Gigaword GPT-3.5 $\ \ -10.37_{\pm 11.34}$
GPT-4 $\ \ -7.55_{\pm 4.54}$
Llama-7B $\ -19.33_{\pm 1.17}$ $5.23$
Mistral-7B $\ -22.56_{\pm 7.32}$ $3.13$
Rotten GPT-3.5 $\ \ \ \ 1.77_{\pm 3.81}$
Tomatoes GPT-4 $\ \ -9.21_{\pm 3.08}$
Llama-7B $\ -13.07_{\pm 4.56}$ 9.91
Mistral-7B $\ \ \ 12.24_{\pm 5.58}$ $0.83$
TweetQA GPT-3.5 $\ \ -14.14_{\pm 11.87}$
GPT-4 $\ \ -7.55_{\pm 2.65}$
Llama-7B $\ -12.28_{\pm 4.44}$ $6.37$
Mistral-7B $\ \ -8.48_{\pm 5.95}$ $2.78$
Table 2: Task-switch impact from CH-tasks ($T_{H}$) to target ($T_{t}$): MMLU AA and conversation length L = 6. Sensitivity not calculable for *.
@@
Rotten GPT-3.5 $\ \ \ 3.00_{\pm 1.41}$
Tomatoes GPT-4 $\ \ \ 1.74_{\pm 1.14}$
Llama-7B $\ \ \ 2.54_{\pm 0.55}$ $4.02$
Mistral-7B $\ \ \ 3.17_{\pm 0.47}$ $2.65$
Gigaword GPT-3.5 $\ \ \ 0.11_{\pm 1.63}$
GPT-4 $\ -0.98_{\pm 2.39}$
Llama-7B $\ \ \ 1.82_{\pm 0.18}$ 1.98
Mistral-7B $\ -0.79_{\pm 0.55}$ 3.04
MMLU AA GPT-3.5 $\ -0.22_{\pm 1.25}$
GPT-4 $\ \ \ 0.76_{\pm 0.69}$
Llama-7B $\ -5.33_{\pm 0.22}$ 3.37
Mistral-7B $\ \ \ 1.33_{\pm 0.23}$ 1.39
TweetQA GPT-3.5 $\ -0.33_{\pm 3.35}$
GPT-4 $\ -0.98_{\pm 1.49}$
Llama-7B $\ \ \ 2.72_{\pm 0.53}$ 2.77
Mistral-7B $\ -1.23_{\pm 0.23}$ 3.01
@@
Table 3: Task-switch impact from CH-tasks ($T_{h}$) to target ($T_{t}$): Rotten Tomatoes and conversation length L = 6. Sensitivity not calculable for *.
@@
In practice, where there is no access to ground-truth reference responses, the performance degradation of deployed models cannot be computed for real conversational task-switches, but instead the sensitivity of different models to different task-switches can be predicted using the reference-free task-switch metric, $\tau(\cdot)$, as introduced in Equation 3. The larger the value of $\tau(\cdot)$, the greater a model’s sensitivity to a specific task-switch. In Table 2 and Table 3, Llama-7B usually has the highest sensitivity to task-switches with for example $\tau=3.37$ for a switch from MMLU AA to Rotten Tomatoes and $\tau=9.91$ for task-switch from Rotten Tomatoes to MMLU AA. We observe a general trend between the change in accuracy and $\tau(\cdot)$ for task-switch scenarios for $T_{t}=$ Rotten Tomatoes where a negative change in performance also suggests very high task-switch sensitivity. In Figure 2, we plot the change in performance with increasing $T_{h}$ examples for MMLU AA dataset. Performance fluctuations for conversation history, $\mathbf{h}$, can stem from two primary factors: a significant drop in the predicted probability for the zero-shot response, $r^{*}$, or a notable increase in the probability for an alternative response, $r$. The latter can result in substantial performance change while maintaining low sensitivity, $\tau(\cdot)$. By analyzing both performance changes and task-switch sensitivity, we gain deeper insights into the models’ adaptability to task-switches and the underlying dynamics influencing these shifts.
4.3 Discussion
Task-switch is demonstrably a threat to model performance in conversational contexts. The reference-free task-switch sensitivity metric, $\tau(\cdot)$ offers a powerful tool to predict the extent to which a model may be vulnerable to an observed task-switch in a real conversation for a deployed model. However, it is further useful to understand why certain models and certain task-switches lead to greater performance degradation than others. In this section, two natural hypotheses are explored. First, it is posited that there may be a relationship between the average length of a response for the conversation-history task, $T_{h}$ and the performance degradation for a particular target task, $T_{t}$. The length metric behaves as a single-value proxy metric for characterizing the difference between different tasks. This is evaluated empirically in Appendix F.1 and no such correlation is found unfortunately. Considering the average task response length can be viewed as an attempt to measure the distance between the conversation-history task, $T_{h}$ and the target task, $T_{t}$. Therefore, a second natural hypothesis is that a more formal distance measure can better explain the extent of performance degradation for different task-switches. Experiments in Appendix F.2 explore a ranking based approach as per a range of powerful Large Language Models (LLMs) to obtain some measure of the distance between tasks. The empirical results show a weak correlation, which perhaps leads to the conclusion that the performance degradation cannot be simply explained by the specific task-switch but is also a function of the specific model. Future work is required to further understand comprehensively the components that explain the variations in performance degradation for different task-switches and different models..
@@ 5 Conclusions and Future Work
This work formalizes and investigates the sensitivity of large language models (LLMs) to task-switch scenarios within conversational contexts. We introduce a task-sensitivity as a reference-free metric that can explain a model’s behavior to task-switches along with the performance change. By experimenting with various task-switch settings, we observe that even advanced models like GPT-4 can exhibit vulnerabilities to task-switches. Our work additionally lays the foundation for future work on ‘side-channel’ vulnerabilities of LLMs to undesired information leakage/bias from the conversation history. Further work will focus on developing adaptive context management strategies within LLMs to mitigate the risk of task-switch sensitivity.
@@ 8 Acknowledgements
Appendix A gives more details about the datasets, Appendix B reports the zero-shot absolute performance of all models on all tasks, Appendix C presents an ablation study on the conversation history length (with multiple seeds), Appendix D discusses the prompt templates, Appendix E discusses other definitions for task-switch sensitivity, Appendix F discusses correlations, Appendix G tabulates confusion matrices for each model, Appendix H investigates task-switch without teacher-forcing the responses in the history, and Appendix I studies the impact of a randomly generated conversation history.
@@ Appendix B Absolute Performance
When evaluating the target task with a conversation history, it is useful to compare the performance against a baseline with no conversation history ($\mathbf{h}=\text{Ø},L=0$). This is equivalent to evaluating in a zero-shot setting. This section reports the zero-shot performance for all the target task ($T_{t}$) datasets: MMLU HA in Table  5, MMLU AA in Table 6, RT in Table 7, Gigaword in Table  8 and TweetQA in Table 9. Also note that for the classification tasks (MMLU HA, MMLU AA, RT), we also report the number of responses for which we were unable to extract the answer (# Format Errors), which is further discussed in Appendix D. We evaluated the test set with four LLMs (GPT-3.5, GPT-4, Mistral-7B, Llama-7B), which were all set to temperature 0 for reproducibility.